OpenStudy (anonymous):

The distance of the school is 3 kilometer how long will kilometers will you travel round trip between Monday and Friday

OpenStudy (jamesj):

@meltedcheese, for the record, it's considered bad form to buzz off half way through someone answering your question.)

OpenStudy (anonymous):

sorry my iPad only has a limit of answering questions... It runs out of spaace

OpenStudy (unklerhaukus):

\[3(km)\times 5 (days/montofri) \times 2 \times (there) per (there and back)\]

OpenStudy (anonymous):

3x5x2

OpenStudy (unklerhaukus):

30 km

OpenStudy (anonymous):

Ohh I get it thanks!